# A is for Absorption: Studying Feature Splitting and Absorption in Sparse Autoencoders

### v5

@@ 2. Proposition 1: Perfect Reconstruction under $\delta$-Absorption
For any $\delta\in[0,1]$, and for inputs $h$ consisting only of $f_{1},f_{2}$ or 00, the reconstruction $\hat{h}=W_{d,1}z_{1}+W_{d,2}z_{2}$ perfectly reconstructs $h$, i.e., the reconstruction loss component $\mathcal{L}_{\text{rec}}^{(1,2)}$ due to these features is 00.
@@
[…]
@@
[…]
@@ 4. Corollary: Increasing Absorption Decreases Sparsity Loss
From Proposition 2, $\frac{d\mathcal{L}_{\text{sp}}^{(1,2)}}{d\delta}=-p_{11}$. If $p_{11}>0$, then $-p_{11}<0$. A negative derivative implies that $\mathcal{L}_{\text{sp}}^{(1,2)}$ is a decreasing function of $\delta$ for $\delta\in[0,1]$. The minimum value of $\mathcal{L}_{\text{sp}}^{(1,2)}$ over this interval occurs at $\delta=1$ (full absorption), yielding $\mathcal{L}_{\text{sp}}^{(1,2)}(\delta=1)=p_{11}(2-1)+p_{10}=p_{11}+p_{10}$. The maximum value occurs at $\delta=0$ (no absorption), yielding $\mathcal{L}_{\text{sp}}^{(1,2)}(\delta=0)=p_{11}(2-0)+p_{10}=2p_{11}+p_{10}$. $\square$
@@ A.3 Extended toy model experiments
Figure 10: SAE encoder and decoder with true features. The firing magnitude of feature 0 has variance 0.1, while the remaining features fire with variance 0.0. When there is variance in the firing magnitudes of parent and child features, we still see an absorption pattern in the SAE encoder and decoder with the latent 3 encoder containing a negative component of feature 1, and the latent 0 decoder merging features 0 and 1.
@@
So far, we have only shown feature absorption occurring with standard L1 SAEs. Next, we examine how other absorption affects other architectures using a batch topk SAE [3]. Batch topk SAEs are an improved version of topk SAEs [10] where the top $k*B$ latents are used to reconstruct the SAE input, where $B$ is the batch size. As the topk function enforces sparsity, there is no additional L1 loss term.
@@ A.4 Ablation algorithm
  Insert SAE in model computation, including error term
  Define a scalar metric on the model’s output distribution
  Calculate baseline metric value for a test prompt
  for each token of interest do
     for each SAE latent do
        Set the SAE latent activation to 0
        Recalculate the metric
        Compute ablation effect (baseline - new metric)
        Reset the SAE latents to its original value
     end for
  end for
@@ A.9 SAE training
We train SAEs on the first 8 layers of Qwen2 0.5B [32] and Llama 3.2 1B [6] using the SAELens library [15]. The SAEs are all trained with identical hyperparameters of L1 coefficient of 2.5 and 500M tokens. The Qwen2 0.5B SAEs all have L0 between 25 and 50 and explained variance between 0.77 and 0.83. The Llama 3.2 1B SAEs have L0 between 27 and 110, and explained variance between 0.74 and 0.89. We use a single 40gb Nvidia A100 GPU for training each SAE.
@@ A.12 Causal interventions and absorption
During a LLM forward pass, the model first collects relevant information on a token in that token position, and attention heads then move relevant information from earlier tokens to later tokens [11, 22]. If we assess ablation effect at layers after which model attention has already pulled relevant information from the subject tokens into the final output token, the ablation effect will be 0. For Gemma 2 2B on the first-letter spelling task, we find this movement of first-letter spelling information occurs around layer 18.
Figure 22 shows an activation patching experiment [22] on a sample first-letter spelling prompt. In this experiment, we see that near layer 18 the model moves first-letter spelling information from the subject token to the prediction token.

### v6

@@ 2. Proposition 1: Perfect Reconstruction under $\delta$-Absorption
For any $\delta\in[0,1]$, and for inputs $h$ consisting only of $f_{1},f_{2}$ or 0, the reconstruction $\hat{h}=W_{d,1}z_{1}+W_{d,2}z_{2}$ perfectly reconstructs $h$, i.e., the reconstruction loss component $\mathcal{L}_{\text{rec}}^{(1,2)}$ due to these features is 0.
@@
$=\text{ReLU}(1-\delta\cdot 0)=1\quad(\text{by H1})$
$z_{2}$ $=\text{ReLU}(W_{e,2}\cdot h)=\text{ReLU}(f_{2}\cdot f_{1})=\text{ReLU}(0)=0\quad(\text{by H1})$
@@
$=\text{ReLU}(1+0-\delta\cdot 0-\delta\cdot 1)=\text{ReLU}(1-\delta)\quad(\text{by H1})$
@@ 4. Corollary: Increasing Absorption Decreases Sparsity Loss
From Proposition 2, $\frac{d\mathcal{L}_{\text{sp}}^{(1,2)}}{d\delta}=-p_{11}$. If $p_{11}>0$, then $p_{11}<0$. A negative derivative implies that $\mathcal{L}_{\text{sp}}^{(1,2)}$ is a decreasing function of $\delta$ for $\delta\in[0,1]$. The minimum value of $\mathcal{L}_{\text{sp}}^{(1,2)}$ over this interval occurs at $\delta=1$ (full absorption), yielding $\mathcal{L}_{\text{sp}}^{(1,2)}(\delta=1)=p_{11}(2-1)+p_{10}=p_{11}+p_{10}$. The maximum value occurs at $\delta=0$ (no absorption), yielding $\mathcal{L}_{\text{sp}}^{(1,2)}(\delta=0)=p_{11}(2-0)+p_{10}=2p_{11}+p_{10}$. $\square$
@@ A.3 Extended toy model experiments
Figure 10: SAE encoder and decoder with true features. The firing magnitude of feature 0 has variance 0.1, while the remaining features fire with variance $0.$ When there is variance in the firing magnitudes of parent and child features, we still see an absorption pattern in the SAE encoder and decoder with the latent 3 encoder containing a negative component of feature 1, and the latent 0 decoder merging features 0 and 1.
@@
So far, we have only shown feature absorption occurring with standard L1 SAEs. Next, we examine how other absorption affects other architectures using a batch topk SAE [bussmann2024batchtopksparseautoencoders]. Batch topk SAEs are an improved version of topk SAEs [gao2024scaling] where the top kBk*B latents are used to reconstruct the SAE input, where $B$ is the batch size. As the topk function enforces sparsity, there is no additional L1 loss term.
@@ A.4 Ablation algorithm
1:Insert SAE in model computation, including error term
2:Define a scalar metric on the model’s output distribution
3:Calculate baseline metric value for a test prompt
4:for each token of interest do
5:  for each SAE latent do
6:   Set the SAE latent activation to 0
7:   Recalculate the metric
8:   Compute ablation effect (baseline - new metric)
9:   Reset the SAE latents to its original value
10:  end for
11:end for
@@ A.9 SAE training
We train SAEs on the first 8 layers of Qwen2 0.5B [qwen2] and Llama 3.2 1B [dubey2024llama3herdmodels] using the SAELens library [bloom2024saetrainingcodebase]. The SAEs are all trained with identical hyperparameters of L1 coefficient of 2.5 and 500M tokens. The Qwen2 0.5B SAEs all have L0 between 25 and 50 and explained variance between 0.77 and 0.83. The Llama 3.2 1B SAEs have L0 between 27 and 110, and explained variance between 0.74 and 0.89. We use a single 40gb Nvidia A100 GPU for training each SAE.
@@ A.12 Causal interventions and absorption
During a LLM forward pass, the model first collects relevant information on a token in that token position, and attention heads then move relevant information from earlier tokens to later tokens [geva2023dissecting, meng2022locating]. If we assess ablation effect at layers after which model attention has already pulled relevant information from the subject tokens into the final output token, the ablation effect will be 0. For Gemma 2 2B on the first-letter spelling task, we find this movement of first-letter spelling information occurs around layer 18.
Figure 22 shows an activation patching experiment [meng2022locating] on a sample first-letter spelling prompt. In this experiment, we see that near layer 18 the model moves first-letter spelling information from the subject token to the prediction token.